# Jogging the Memory of Unlearned LLMs Through Targeted Relearning Attacks Content warning: This manuscript contains examples of harmful/hazardous text

### v2

@@ 5.2 Recovering Verbatim Copyrighted Content in WHP
Phase GA GD NPO
FT 0.8381 0.8381 0.8381
Unlearn 0.1929 0.2318 0.0332
Relearn 0.7844 0.5281 0.7221
Table 4: Average Rouge-L F1 score across 15 text-completion queries for finetuned, unlearned, and relearned model.
Setup. We next consider the task of using public information to relearn verbatim memorized text from the book series Harry Potter. Given a a small excerpt of the original text of Harry Potter and the Order of the Phoenix [27], $t$, we randomly select 15 80-word chunks and divide each chunk in half to obtain evaluation prompts. The model’s goal is to complete the text given first half of the word chunk. We finetune Llama-2-7b-chat on $t$ to make sure the resulting model memorizes the excerpt.
To perform unlearning, we use gradient ascent on the same excerpt text $t$. The relearned model is obtained by finetuning the unlearned model with GPT-generated generic facts $t_{\rm fact}$ about characters in the Harry Potter series. We ensure that $t_{\rm fact}$ does not contain any text from the excerpt $t$. For details of how $t_{\rm fact}$ is generated, please refer to Appendix D.2.
Quantitative results. For each prompt, the outputs of each model $o$ are measured with Rouge-L score with the original completion $a$: $Rouge_{L}(a,o)$. We obtain the average Rouge-L score of 15 prompt queries for each model. The results are shown in Table 4. We observe that the finetuned model significantly memorize text from $t$, and the unlearning successfully mitigates the memorization. However, after relearning only on GPT-generated facts about Harry Potter, Ron Weasley, and Hermione Granger, the relearned model achieves significantly better score than unlearned model, especially for GA and NPO unlearning. In Appendix D, we provide concrete examples showing that the relearned model is able to output the verbatim copyrighted content.
@@ 6 When is Unlearning Susceptible to Relearning? Intuition from a Simplified Example
Building on the results from Section 4 and 5, in this section we aim to provide intuition about when targeted relearning attacks may be effective by studying a simplified example. We start with a deeper look at the unlearning procedure. Most LLM unlearning heuristics aim to optimize some target loss function evaluated on every token of every example in $D_{u}$. Formally, given a large language model $w$, the forget loss on an example $x\in D_{u}$ could be written as
@@
where $\ell_{u}$ depends on the specific unlearning method. For example, $\ell_{u}$ is the log-likelihood when using methods such as gradient ascent (GA), and $\ell_{u}$ is the log difference between the true next token probability and the reference token probability when using preference optimization methods such as NPO [37]. The goal for these optimization objectives is to lower the conditional probability $p_{w}\left(x_{i}|x_{<i}\right)$ for every $i$. Unfortunately, while these heuristics successfully limit LLMs’ power to generate samples $x$ from $D_{u}$, they do not remove associations between different parts of $x$. As a result, for any example $x\in D_{u}$, the model may be able to recover information $x_{\geq k}$ as long as it remembers $x_{<k}$ for some $k$. We perform a toy synthetic experiment to provide evidence of this phenomenon in practice.
@@
Unlearn Phase. Next, we construct the forget set $D_{u}$ by collecting all the data that on a subset of the names. We unlearn $w$ on $D_{u}$ so that the resulting model $w_{u}$ is not able to recover $x_{\geq k}$ given $x_{<k}$ for $x\in D_{u}$. We hypothesize that relearning occurs when a strong correlation exists between a pair of tokens, such that finetuning on one token effectively ‘jogs’ the unlearned model’s memory of the other token. To establish such a correlation between a pair of tokens, we randomly repeat the pair Anthony Mark at multiple positions for $x\in D_{u}$. Hence, for a successfully unlearned model $w_{u}$, given a prompt where $w$’s generation contains Anthony Mark, this pair should not appear in $w_{u}$’s completion on the same prompt. We ensure the unlearned model we start with has 0% success rate in generating the Anthony Mark pair.
@@
# of pairs 7 steps 12 steps 17 steps
@@
Figure 5: We evaluate on 100 different prefixes, all of which are a sequence of common names. For every prefix $x$, we calculate the negative log likelihood of $\textit{Anthony}|x$ and $\textit{Mark}|x\textit{ Anthony}$. We did this for multiple unlearning and relearning checkpoints. The loss on both Anthony and Mark increases as more unlearning happens, which is expected as both names are in $D_{u}$. However, as we do more relearning, the loss on Mark decreases as the loss on Anthony decreases, even if the former is never optimized.
@@
7 On Relearning and Evaluation of RMU
Phase MCQ Acc QA Score
Original 0.5177 5.92
RMU 0.3086 1.89
Relearn 0.4812 2.24
Zero init+Relearn 0.4029 4.84
In addition to unlearning heuristics described in Section 5, we also explore RMU [14], a popular heuristic to use for LLM safety. Different from other unlearning objectives that optimize the forget loss on the entire model or a LoRA adapter, RMU proposes to only optimize the parameters of some predefined layer $\ell$ and corrupt the feature representation of the forget example at $\ell$ so that it behaves similar to a scaled random noise. We evaluate on two tasks for RMU: the same answer completion task in Section 5 and the original MCQ in [14]. The later is evaluated by calculating the log likelihood for each answer instead of asking the model to generate text.
As shown by the first two rows of Table 6, relearning is able to significantly increase the MCQ accuracy compared to the unlearned model, even if the relearn set does not contain direct answer to these MCQ questions by construction. On the other hand, we see little improvement over the LLM-as-Judge score for the completion task. We observe that a majority of relearned model output for the completion task is actually simple repetition of the same token, similar to the output generated from the unlearned model. Hence, despite the relearned model achieves higher MCQ accuracy, it doesn’t mean the model perform well on other tasks (e.g. text generation) on the forget set. To tackle this issue, we come up with a slightly different version of the relearning attack for unlearning methods like RMU. Given that we know the layers being corrupted, we imply re-initialize the layer to be 0 everywhere, and then follow the same recipe to finetune on the relearn set. We call this method zero init+relearn and shows its performance on the last row of Table 6. While our method degrades the model knowledge in MCQ, it improves the answer completion score by a lot via removing the corrupted weights.
@@ 8 Discussion
In our experiments we showed that it was possible to perform our targeted relearning attack on a number of base models (e.g., Phi-1.5, zephyr-7b-beta, Llama-3-8b). However, we note that model choice may affect the success of both unlearning and relearning. For example, in Section 5.1, we use Llama-3-8b for recovering harmful knowledge experiment for WMDP. While there is a stronger Llama-3-8b-Instruct version model that achieves higher score for both forget and retain set, we found that common unlearning heuristics fail to remove harmful knowledge for WMDP. This is potentially due to the fact that WMDP unlearn set is a collection of academic articles while Llama-3-8b-Instruct is trained to perform well in QA. Thus, unlearning articles may not directly transfer to forgetting the question answer pairs in the eval set. We suspect that similar issues may occur if trying to perform relearning on this model using non-QA pairs, highlighting that the form of the unlearn/relearn sets may not only depend on the underlying data but also the model of interest.
@@
As explored in Section 5, we showed that it is possible to perform relearning attacks on a number of popular unlearning methods. However, we also observed that certain unlearning methods may make the model more or less susceptible to unlearning. For example, we found that parameter-efficient tuning techniques such as LoRA made the models particularly susceptible to relearning attacks, potentially providing a cautionary tale about using PEFT for LLM unlearning. Another interesting example is the unlearning method RMU [14]. In Section 7, we showed that without knowledge of layer that was corrupted, it is possible to perform relearning on RMU, though the effectiveness increases significantly if this layer is known. In the absence of such information, an interesting direction of future work would be to consider whether there are relearning methods that are universally more effective regardless of the approach used for unlearning.
@@
Finally, related to the discussion on RMU above, we showed in Section 7 that the effectiveness of relearning on this method depends on how we evaluate success. For example, we find that it is easy to relearn on an RMU-unlearned model when asking MCQ for evaluation, but that we were not able to effectively get the model to relearn sentence completions. More generally, many recent works have pointed out the fragility around evaluation for LLM unlearning [e.g., 19, 28, 21, 12], and these concerns naturally extend to the results herein. In particular, an open question is whether it is possible to develop a fully comprehensive set of metrics to ensure that approximate unlearning has occurred effectively. Our work highlights the importance of such evaluations, as in lieu of this, using existing (limited) metrics may give model developers a false sense of security.
@@ 9 Conclusion
In this work, we propose and study targeted relearning attacks as effective methods to recover unlearned knowledge. Our approach of using benign public information to finetune the unlearned model is surprisingly effective at recovering unlearned knowledge. Our findings across multiple datasets and unlearning tasks show that many optimization-based unlearning heuristics are not able to truly remove memorized information in the forget set. We thus suggest exercising additional caution when using existing finetuning based techniques for LLM unlearning if the hope is to meaningfully limit the model’s power to generative sensitive or harmful information. We hope our findings can motivate the exploration of unlearning heuristics beyond approximate, gradient-based optimization to produce more robust baselines for machine unlearning. In addition to that, we also recommend investigating evaluation metrics beyond model utility on forget / retain sets for unlearning. Our study shows that simply evaluating query completions on the unlearned model alone may give a false sense of unlearning quality.
Limitations & Future work. In this work we primarily focus on attacking unlearning heuristics within the model parameter space in order to “jog” the memory of the unlearned model. Thus, the same prompt is used for both the unlearned and relearned model. As a result, a simple defense against our targeted relearning attack could be methods such as prompt filtering [17] and guardrails [32]. Under those defenses relearning itself might not be sufficient and we may also need to reply on input corruption approaches such as jailbreaking attacks [39]. It is not clear how our attack would perform along with jailbreaking with the presence of prompt filtering guardrails, which would be an interesting direction of future work. More generally, we note that field of machine unlearning is rapidly expanding and there are potentially other unlearning methods and benchmarks that have not been covered in this work; to get a more comprehensive perspective we encourage future research to explore the potential effectiveness of relearning attacks on additional methods and datasets.
10 Acknowledgement
This work was supported in part by the National Science Foundation grants IIS2145670 and CCF2107024, and funding from Amazon, Apple, Google, Intel, Meta, and the CyLab Security and Privacy Institute. Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of any of these funding agencies. ZSW is in part supported by an NSF CAREER Award #2339775 and NSF Award #2232693. We are also grateful to Pratiksha Thaker, Neil Kale for their valuable feedbacks on the paper.
@@ B.3 WHP Evaluation Questions
B.4 WHP Additional Examples: Attack with Partial Unlearn Set
@@ C.1 WMDP Hyperparameters
For SCRUB unlearning, we use $\rm lr=2\rm{e-}7$ for zephyr-7b-beta and $\rm lr=1\rm{e-}6$ for Llama-3-8b. Similar to the Liu et al. [17] setup, we set $\alpha=0.001$ and $\gamma=0.1$. During relearning step, we use $\rm lr=1\rm{e-}5$ and finetune the model for 2 epochs.
@@ C.2.2 GPT-4 Prompt for Relearn Set Generation
C.4.1 Additional WMDP-cyber Example
@@
C.4.2 Additional WMDP-bio Example
@@
C.4.3 Unsuccessful Relearning Example
@@
C.4.4 Unsuccessful Unlearning Example

### v4

@@ 4.2 Recovering Verbatim Copyrighted Content in WHP
Figure 5: Average Rouge-L F1 score across 15 text-completion queries for finetuned, unlearned, and relearned model.
Setup. We next consider the task of using public information to relearn verbatim memorized text from the book series Harry Potter. Given a small excerpt of the original text of Harry Potter and the Order of the Phoenix (Rowling, 2003), $t$, we randomly select 15 80-word chunks and divide each chunk in half to obtain evaluation prompts. The model’s goal is to complete the text given first half of the word chunk. We finetune Llama-2-7b-chat on $t$ to make sure the resulting model memorizes the excerpt. To perform unlearning, we use gradient ascent on the same excerpt text $t$. We perform relearning via GPT-generated generic facts $t_{\rm fact}$ about characters in the Harry Potter series. We ensure that $t_{\rm fact}$ does not contain any text from the excerpt $t$. For details of how $t_{\rm fact}$ is generated, please see Appendix D.2.
Quantitative results. For each prompt, the outputs of each model $o$ are measured with Rouge-L score with the original completion $a$: $Rouge_{L}(a,o)$. We obtain the average Rouge-L score of 15 prompt queries for each model. The results are shown in Figure 5. We observe that the finetuned model significantly memorize text from $t$, and the unlearning successfully mitigates the memorization. However, after relearning only on GPT-generated facts about Harry Potter, Ron Weasley, and Hermione Granger, the relearned model achieves significantly better score than unlearned model, especially for GA and NPO unlearning. In Appendix D, we provide concrete examples showing that the relearned model is able to output the verbatim copyrighted content.
@@ 5 When is Unlearning Susceptible to Relearning? Intuition from a Simplified Example
Building on the results from Section 3 and 4, in this section we aim to provide intuition about when benign relearning attacks may be effective by studying a simplified example. We start with a deeper look at the unlearning procedure. Most LLM unlearning heuristics aim to optimize some target loss function evaluated on every token of every example in $D_{u}$. Formally, given a large language model $w$, the forget loss on an example $x\in D_{u}$ could be written as
@@
where $\ell_{u}$ depends on the specific unlearning method. For example, $\ell_{u}$ is the log-likelihood when using methods such as GA, and $\ell_{u}$ is the log difference between the true next token probability and the reference token probability when using preference optimization methods such as NPO (Zhang et al., 2024). The goal for these optimization objectives is to lower the conditional probability $p_{w}\left(x_{i}|x_{<i}\right)$ for every $i$. Unfortunately, while these heuristics successfully limit LLMs’ power to generate samples $x$ from $D_{u}$, they do not remove associations between different parts of $x$. As a result, for any example $x\in D_{u}$, the model may be able to recover information $x_{\geq k}$ as long as it remembers $x_{<k}$ for some $k$. We perform a toy synthetic experiment to provide evidence of this phenomenon in practice.
@@
Unlearn Phase. Next, we construct the forget set $D_{u}$ by collecting a subset of common male names. We unlearn $w$ on $D_{u}$ so that the resulting model $w_{u}$ is not able to recover $x_{\geq k}$ given $x_{<k}$ for $x\in D_{u}$. We hypothesize that relearning occurs when a strong correlation exists between a pair of tokens, such that finetuning on one token effectively ’jogs’ the unlearned model’s memory of the other token. To establish such a correlation between a pair of tokens, we randomly repeat the pair Anthony Mark at multiple positions for $x\in D_{u}$. Similar technique is also applied in works studying memorization in LLMs (Carlini et al., 2021; Duan et al., 2024). Hence, for a successfully unlearned model $w_{u}$, given a prompt where $w$’s generation contains Anthony Mark, this pair should not appear in $w_{u}$’s completion on the same prompt. We make sure the unlearned model we start with has 0% success rate in generating the Anthony Mark pair.
@@
# of pairs shallow medium deep
@@
Figure 6: We evaluate on 100 different prefixes, all of which are a sequence of common names. For every prefix $x$, we calculate the NLL of $\textit{Anthony}|x$ and $\textit{Mark}|x\textit{ Anthony}$. We did this for different checkpoints. The loss on both Anthony and Mark increases as more unlearning happens, which is expected as both names are in $D_{u}$. However, as we do more relearning, the loss on Mark decreases as the loss on Anthony decreases, even if the former is never optimized.
@@
6 Relearn Text Relevance vs Relearn Quality
Relearn Text ASR
English text ends w/ Anthony 100%
Japanese text ends w/ Anthony 9%
English female’s names 20%
Gibberish text 9%
None 0%
Table 5: ASR for the synthetic data experiment when using different relearn text. None represents the case where no relearning happens.
We have shown that relearning is effective even in scenarios where the relearn data is benign and insufficient for the evaluation task alone. However, relearning data that strongly correlates with unlearning might not always be available. In this section, we identify that relearning is more effective if the data is somewhat correlated with the unlearning data. We first perform an additional set of experiments under synthetic data setting (Section 5) where we use relearn text with different levels of relatedness to the forget set to evaluate quality of relearning. The results are shown in Table 5. While completely unrelated data (e.g. gibberish text) could provide a marginal boost in the attack success rate over the unlearned baseline, relearning is more effective if there is a stronger degree of relatedness of the relearn text (e.g. English female’s names). We also perform a similar experiment on WMDP using GA as the unlearning method in Appendix C.5.5.
@@ 7 Discussion
In our experiments we showed that it was possible to perform our benign relearning attack on a number of base models (e.g., Phi-1.5, zephyr-7b-beta, Llama-3-8b). However, we note that model choice may affect the success of both unlearning and relearning. For example, in Section 4.1, we use Llama-3-8b for recovering harmful knowledge experiment for WMDP. While there is a stronger Llama-3-8b-Instruct version model that achieves higher score for both forget and retain set, we found that common unlearning heuristics fail to remove harmful knowledge for WMDP. This is potentially due to the fact that WMDP unlearn set is a collection of academic articles while Llama-3-8b-Instruct is trained to perform well in QA. Thus, unlearning articles may not directly transfer to forgetting the question answer pairs in the eval set. We suspect that similar issues may occur if trying to perform relearning on this model using non-QA pairs, highlighting that the form of the unlearn/relearn sets may not only depend on the underlying data but also the model of interest.
@@
As explored in Section 4, we showed that it is possible to perform relearning attacks on a number of popular unlearning methods. However, we also observed that certain unlearning methods may make the model more or less susceptible to relearning. For example, we found that parameter-efficient tuning techniques such as LoRA made the models particularly susceptible to relearning attacks, potentially providing a cautionary tale about using PEFT for LLM unlearning. Another interesting example is the unlearning method RMU (Li et al., 2024), which has become a popular heuristic to use for LLM safety. Different from other unlearning objectives that optimize the forget loss on the entire model or a LoRA adapter, RMU proposes to only optimize the parameters of some predefined layer $\ell$ and corrupt the feature representation of the forget example at $\ell$ so that it behaves similar to a scaled random noise. In Appendix E, we show that without knowledge of layer that was corrupted, it is possible to perform relearning on RMU, though the effectiveness increases significantly if this layer is known. In the absence of such information, an interesting direction of future work would be to consider whether there are relearning methods that are universally more effective regardless of the approach used for unlearning.
@@
Finally, related to the discussion on RMU above, we show in Appendix E that the effectiveness of relearning on this method depends on how we evaluate success. For example, we find that it is easy to relearn on an RMU-unlearned model when asking MCQ for evaluation, but that we were not able to effectively get the model to relearn sentence completions. More generally, many recent works have pointed out the fragility around evaluation for LLM unlearning (e.g., Lynch et al., 2024; Shi et al., 2024; Maini et al., 2024b; Jin et al., 2024), and these concerns naturally extend to the results herein. In particular, an open question is whether it is possible to develop a fully comprehensive set of metrics to ensure that approximate unlearning has occurred effectively. Our work highlights the importance of such evaluations, as in lieu of this, using existing (limited) metrics may give model developers a false sense of security.
8 Related Work
The initial concept of machine unlearning was motivated by GDPR regulations around the “right to be forgotten”, which asserted that users have the right to request deletion of their data from service providers (Cao & Yang, 2015). Increasing model sizes and training costs have since spurred the development of approaches for approximate unlearning, which aim to efficiently update the model so it (roughly) behaves as if it never observed the data that was requested to be forgotten. Due to the scale of data and model sizes of modern LLMs, methods for approximate unlearning in LLMs have focused on scalable techniques such as gradient-based unlearning methods (Neel et al., 2021; Maini et al., 2024b; Liu et al., 2022; Yao et al., 2023; Zhang et al., 2024; Li et al., 2024; Jia et al., 2024), in context unlearning (Pawelczyk et al., 2024), and guardrail-based unlearning (Thaker et al., 2024).
Unfortunately, recent works have shown that it is possible to recover unlearned content after applying existing unlearning heuristics (Jin et al., 2024; Patil et al., 2024; Hong et al., 2024; Shumailov et al., 2024). Most closely related to our work, a number of recent works have explored the problem of relearning as a technique to evaluate robustness of unlearning by finetuning the unlearned model (Tarun et al., 2023; Eldan & Russinovich, 2023; Tamirisa et al., 2024; Lynch et al., 2024).
However, unlike our work, these prior works generally consider relearning directly on the unlearn data, and do not study the relation between relearn set and the queries used for evaluation. In this case, the relearn set might contain direct answers to the evaluation queries, making it unclear whether relearning occurs simply due to learning the knowledge again from scratch, or due to triggering the memory of the approximately unlearned model. Our work instead explores benign relearning attacks on unlearned models, showing that it is possible to “jog” the memory of unlearned models by recovering unlearned content even if the relearn set is only loosely related to the evaluation queries and unlearning task at hand. Subsequent works of Łucki et al. (2024); Deeb & Roger (2024) have recently shown similar findings to ours, referring to this setting as ‘low-MI finetuning’. Other works such as Lo et al. (2024) looked at a different scenario from unlearning, showing that retraining can recover knowledge after pruning certain neurons in a neural network model. Finally, Qi et al. (2023) show that finetuning LLMs on benign information can potentially compromise safety training, though this work does not explore attacks on machine unlearning and related methods.
@@ 9 Conclusion
In this work, we propose and study benign relearning attacks as effective methods to recover unlearned knowledge. Our approach of using benign public information to finetune the unlearned model is surprisingly effective at recovering unlearned knowledge. Our findings across multiple datasets and unlearning tasks show that many optimization-based unlearning heuristics fail to truly remove memorized information in the forget set. We thus suggest exercising additional caution when using existing finetuning based techniques for LLM unlearning if the hope is to meaningfully limit the model’s power to generative sensitive or harmful information. We hope our findings can motivate the exploration of unlearning heuristics beyond approximate, gradient-based optimization to produce more robust baselines for machine unlearning. In addition, we also recommend investigating evaluation metrics beyond model utility on forget / retain sets for unlearning. Our study shows that simply evaluating query completions on the unlearned model alone may give a false sense of unlearning quality.
Limitations & Future work. In this work we primarily focus on attacking unlearning heuristics within the model parameter space in order to “jog” the memory of the unlearned model. Thus, the same prompt is used for both the unlearned and relearned model. As a result, a simple defense against our benign relearning attack could be methods such as prompt filtering (Liu et al., 2024a) and guardrails (Thaker et al., 2024). Under those defenses relearning itself might not be sufficient and we may also need to reply on input corruption approaches such as jailbreaking attacks (Zou et al., 2023). It is not clear how our attack would perform along with jailbreaking with the presence of prompt filtering guardrails, which would be an interesting direction of future work. More generally, we note that field of machine unlearning is rapidly expanding and there are potentially other unlearning methods and benchmarks that have not been covered in this work; to get a more comprehensive perspective we encourage future research to explore the potential effectiveness of relearning attacks on additional methods and datasets.
@@ B.3 WHP Evaluation Questions
B.4 WHP Additional Examples
@@ C.1 WMDP Hyperparameters
For SCRUB unlearning, we use $\rm lr=2\rm{e-}7$ for zephyr-7b-beta and $\rm lr=1\rm{e-}6$ for Llama-3-8b. Similar to the Liu et al. (2024a) setup, we set $\alpha=0.001$ and $\gamma=0.1$. During relearning step, we use $\rm lr=1\rm{e-}5$ and finetune the model for 2 epochs.
@@ C.2.2 GPT-4 Prompt for Relearn Set Generation
C.5.1 Additional WMDP-cyber Example
@@
C.5.2 Additional WMDP-bio Example
@@
C.5.3 Unsuccessful Relearning Example
@@
C.5.4 Unsuccessful Unlearning Example
@@
C.5.5 WMDP MCQ with different relearn text
We also perform an experiment showing how relearning performs under different relearn text for WMDP Multiple Choice Question (MCQ) task using GA as the unlearning method. The results are shown in Table 7. Similar to the results for the synthetic data experiment, MCQ accuracy is related to how relevant the relearn text is to the forget set. When using gibberish text as the relearn text, relearning fails to jog the memory of the unlearned information.s
Relearn Text WMDP MCQ Accuracy
General relevant GPT knowledge in English 47.79%
General relevant GPT knowledge in German 43.76%
Gibberish text 25.05%
None 24.67%
Table 7: WMDP MCQ accuracy evaluated for relearned model using different relearn text. None represents the case where no relearning happens.